# TriForce: Lossless Acceleration of Long Sequence Generation with Hierarchical Speculative Decoding

### v2

@@ Abstract
With large language models (LLMs) widely deployed in long content generation recently, there has emerged an increasing demand for efficient long-sequence inference support. However, key-value (KV) cache, which is stored to avoid re-computation, has emerged as a critical bottleneck by growing linearly in size with the sequence length. Due to the auto-regressive nature of LLMs, the entire KV cache will be loaded for every generated token, resulting in low utilization of computational cores and high latency. While various compression methods for KV cache have been proposed to alleviate this issue, they suffer from degradation in generation quality. We introduce TriForce, a hierarchical speculative decoding system that is scalable to long sequence generation. This approach leverages the original model weights and dynamic sparse KV cache via retrieval as a draft model, which serves as an intermediate layer in the hierarchy and is further speculated by a smaller model to reduce its drafting latency. TriForce not only facilitates impressive speedups for Llama2-7B-128K, achieving up to 2.31$\times$ on an A100 GPU but also showcases scalability in handling even longer contexts. For the offloading setting on two RTX 4090 GPUs, TriForce achieves 0.108s/token—only half as slow as the auto-regressive baseline on an A100, which attains 7.78$\times$ on our optimized offloading system. Additionally, TriForce performs 4.86$\times$ than DeepSpeed-Zero-Inference [2] on a single RTX 4090 GPU. TriForce’s robustness is highlighted by its consistently outstanding performance across various temperatures. The code is available at https://github.com/Infini-AI-Lab/TriForce.
@@ 1 Introduction
Large language models (LLMs) with long-context capability, such as GPT-4 [1], Gemini [42], and LWM [24] continue to emerge and gain proficient application in scenarios including chatbots, vision generation, and financial analysis [43, 9, 54, 35]. However, losslessly serving these LLMs efficiently is challenging. Because of the auto-regressive nature of LLMs, the entire key-value (KV) cache, which stores intermediate key-value states from previous contexts to avoid re-computation, together with model parameters will be loaded into GPU SRAM for every token generated, resulting in low utilization of computational cores. In addition to the large volume of model parameters, the memory footprint of KV cache, which grows linearly with sequence length [33], is emerging as a new bottleneck for long sequence generation.
@@
Recent methodologies have proposed KV cache eviction strategies [47, 53, 28, 17, 13] to mitigate the substantial memory footprint of KV cache, which selectively discard KV pairs from the cache based on a designed eviction policy, allowing models to generate texts with a limited KV cache budget. However, considering that discarded KV pairs cannot be restored and the difficulty in precisely foreseeing which KV pairs will be crucial for future text generation, they struggle with potential information loss, including hallucination and contextual incoherency [49], particularly in long contexts. Such challenges prevent these approaches from boosting speed without sacrificing the performance of models, as illustrated in Figure 1.
@@
Concurrently, speculative decoding, which leverages a lightweight draft model to sequentially predict the next few tokens and let the target model verify the predicted tokens in parallel, is introduced to accelerate LLM inference while provably precisely preserving model output [22, 6, 45]. Nonetheless, deploying it for long sequence generation faces several challenges. First, training draft models to match the context length of target LLMs requires massive computation and it remains questionable whether these small models can achieve the same accuracy with a context length around 1M [4, 32, 48]. Second, we found that draft models with existing training-free methods (e.g., KV cache eviction strategies) can result in poor speculating performance. A continuously increasing divergence [22] is witnessed as the sequence length increases, as shown in Figure 2(a).
@@
Figure 1: Left: TriForce employs retrieval-based drafting and hierarchical speculation to effectively address the dual bottlenecks. It integrates two models and three caches, comprising a draft model, a target model, and a StreamingLLM cache for the draft model, alongside a retrieval cache and a full cache for the target model. The process initiates by repeatedly drafting for $\gamma_{1}$ steps, assisting the target model with retrieved partial KV cache in generating over $\gamma_{2}$ tokens, which will be further verified by the target model using full KV cache. Right: Evaluating the Llama2-7B-128K on a needle retrieval task indicates that KV cache eviction-based methods, such as StreamingLLM, require a trade-off between latency and accuracy. In contrast, our TriForce successfully maintains low latency without sacrificing accuracy.
@@
In pursuit of lossless acceleration, we utilize the lossless feature of speculative decoding as the foundation of our system. An ideal speculative decoding algorithm should (i) be training-free, (ii) maintain a high acceptance rate with long contexts, and (iii) have low-cost drafting. However, two technical challenges need to be addressed to achieve the goal. First, it is not immediately apparent what we can use for low-latency drafting without training a smaller draft model to match the long context length. Second, the key factors for attaining a high acceptance rate with long contexts remain unclear.
@@
Building on these insights, we introduce a hierarchical speculation approach. For a long-context target model (e.g., Llama2-7B-128K [32]), we leverage the original model weights but only with a small proportion (e.g., $3\%$) of KV cache as a draft to tackle the bottleneck of KV cache. Hierarchically, the draft model is further speculated by a lightweight model (e.g., Llama-68M) with StreamingLLM cache to address the bottleneck of model weights. We present TriForce, depicted in Figure 1, a scalable and robust speculative decoding system that integrates retrieval-based drafting and hierarchical speculation, optimized for both on-chip and offloading scenarios. Specifically,
  • In Section 4.1, by maintaining the full cache, we gain the flexibility to select KV pairs, allowing us to devise a superior KV cache selection method, termed retrieval-based drafting. This strategy retrieves required context information for future needs, which is characterized as lossless, particularly in comparison to eviction-based methods such as StreamingLLM and H2O. We further demonstrated its effectiveness and robustness on different datasets.
@@
Figure 2: (a) A continuously increasing natural divergence [22] between the draft model with StreamingLLM or Sliding-window with Re-computation and Llama2-7B-128K is witnessed as the sequence length increases, indicating a falling acceptance rate for speculative decoding with longer contexts. Additionally, temperature sensitivity signals a lack of robustness. (b) Compared with model weights, KV cache gradually becomes another bottleneck with long contexts. (c) KV cache occupies most of the memory as the context length increases.
Empirically, in Section 5, we perform extensive experiments and ablation studies to demonstrate the effectiveness of TriForce. We show that TriForce achieves up to 2.31$\times$ speedup for Llama2-7B-128K on a single A100 GPU on top of Hugging Face [44] with CUDA graphs [30]. For the offloading setting, TriForce attains an impressive 7.78$\times$ on two RTX 4090 GPUs, reaching 0.108 s/token—only half as slow as the auto-regressive baseline on an A100. TriForce can efficiently serve a Llama2-13B with 128K contexts with 0.226s/token, which is 7.94$\times$ faster on our optimized offloading system. On a single RTX 4090 GPU, TriForce is 4.86$\times$ faster than DeepSpeed-Zero-Inference [2]  [^1] The official implementation of DeepSpeed-ZeRO-Inference with KV cache offloading currently only supports a single GPU, which computes attention on CPU. Our offloading system transfers KV cache from CPU to GPU, benefiting from Tensor Parallelism.. Further, we show that: (i) TriForce has a theoretical 13.1$\times$ upper bound, demonstrating exceptional scalability when dealing with long contexts; (ii) TriForce is robust across various temperature settings, maintaining an acceptance rate above 0.9 even for a temperature of 1.0; and (iii) TriForce’s ability to efficiently process large batches, consistently outperforming the small model with StreamingLLM cache across all configurations, achieving a 1.9$\times$ speedup for a batch size of six, where each sample consists of 19K contexts.
@@ 2.1 Speculative Decoding
Speculative decoding [40, 22, 6, 19, 51, 36, 15] is featured by accelerating LLM decoding while precisely maintaining the model’s output distribution. As the speed of the auto-regressive decoding process is mainly bound by the time for loading model weights and KV cache to GPU SRAM, speculative decoding leverages the observation that generating one token takes the same time as processing tens of tokens in parallel. Tree-based speculation methods are proposed to fully utilize the speculation budget [12, 23]. Instead of making one prediction for the next token, tree-based methods leverage multiple candidates to boost the acceptance rate so that more tokens can get accepted [29, 41, 7]. Staged speculation techniques [39, 8] have been suggested to further accelerate inference by using a cascade of draft models. Meanwhile, self-speculation approaches such as Medusa [5, 3], which are orthogonal to our method, require training efforts and can be integrated into our intermediate draft model.
@@ 2.2 KV Cache Eviction Strategies
StreamingLLM [47] addresses the limitations of window attention and sliding window with re-computation by presenting a straightforward yet effective method that allows LLMs to handle infinitely long text sequences without fine-tuning. StreamingLLM stabilizes the performance by retaining critical attention sink tokens together with recent KV for attention computation. By prioritizing sink tokens, StreamingLLM ensures the attention score distribution remains stable, promoting consistent language modeling for long texts.
H2O [53] introduces a greedy but low-cost approach to processing infinite-length input streams, inspired by a simplified version of the heavy-hitters (H2) eviction policy. This method dynamically updates the KV cache based on the cumulative attention scores, systematically removing the least critical KV to maintain a fixed cache size. By leveraging a greedy algorithm based on local statistics, H2O effectively selects which KV pairs to preserve in the cache, ensuring efficient inference without compromising quality.
However, it is important to recognize that these techniques do not increase the context window size [52, 18, 13, 17]. They focus on retaining only the most recent tokens along with either attention sinks or heavy-hitters, while discarding other tokens. These approaches limit the model to processing based on their designed eviction policies and recent tokens. Consequently, they might not be directly applicable to tasks that demand comprehensive, long-context understanding.
@@ 2.3 KV Cache Quantization
Several approaches to KV cache quantization have been introduced to enhance the efficiency of inference for long sequence generation, aiming to maintain generation quality while reducing the memory consumption [46, 16, 37, 26, 55, 50]. Quantization methods focus on compressing the bit width of KV cache activations, which is orthogonal to our approach.
@@ 3.1 Leveraging Attention Sparsity for Speculative Decoding
The phenomenon of attention sparsity in pre-trained LLMs has been discovered by numerous studies [53, 47, 27, 28]. In our study, we conduct zero-shot inference on the PG-19 test set [34] with Llama2-7B-128K model. By visualizing the sparsity across different attention heads, demonstrated in Figure 3(a), we observe that with a context length of 120K, it is possible to recover over 96% of the attention score with merely 4K tokens across almost all layers.
@@ 4.1 Retrieval-based Drafting
In scenarios requiring long-term contextual dependencies, methods like StreamingLLM and H2O underperform due to their cache updating strategies, which are ineffective at accurately retrieving detailed contextual information because they inevitably and irrecoverably discard KV pairs. In our experiment, we challenge StreamingLLM and H2O with a needle retrieval task [25, 32, 24]. As detailed in Table 1, there is a notable drop in their acceptance rates compared to their performance on the PG-19 dataset, highlighting their limitations. Essentially, StreamingLLM and H2O operate on a lossy principle, as evicted tokens are permanently discarded, making them a poor fit for settings requiring the preservation of full KV cache for the target model.
The necessity of keeping the entire KV cache in our settings allows us to select KV cache more freely [38]. This insight leads us to develop a more effective selection policy for lossless approximations. In our approach, demonstrated in Figure 4, KV cache is segmented into small chunks. During the retrieval phase, we calculate the attention between a given query and the average key cache within each chunk. This method effectively highlights the most relevant chunks, enabling us to gather KV cache with a fixed budget based on the scores. As illustrated in Table 1, retrieval-based method excels by actively identifying the most crucial information for the task rather than relying on passive and time-based cache management methods. By focusing on relevance over recency, retrieval-based policy demonstrates its potential to handle contextually dense datasets.
Table 1: Acceptance rates are shown across various tasks, utilizing a 120K context and a 4K budget, while bypassing the initial two layers. There is a notable drop in StreamingLLM and H2O’s acceptance rate on the needle retrieval task. For reference, Top-K is the theoretical upper bound.
@@ 4.2 Hierarchical Speculation
While addressing the KV cache bottleneck enhances efficiency, the requirement to load whole model weights for drafting reintroduces latency, shifting the bottleneck to model weights again. To tackle this challenge, we implement a hierarchical system, as illustrated in Figure 1. This system employs a secondary, lightweight model with StreamingLLM cache to perform initial speculations for the target model with retrieval-based draft cache (which serves as a draft model for the target model with full KV cache). By establishing this sequential speculation hierarchy, we effectively reduce the drafting latency, thereby accelerating the overall inference.
Correctness: The original output distribution is preserved during the final speculation phase, which is identical to the standard speculative decoding algorithm [22, 6], and the proof is trivial.

### v3

@@ Abstract
With large language models (LLMs) widely deployed in long content generation recently, there has emerged an increasing demand for efficient long-sequence inference support. However, key-value (KV) cache, which is stored to avoid re-computation, has emerged as a critical bottleneck by growing linearly in size with the sequence length. Due to the auto-regressive nature of LLMs, the entire KV cache will be loaded for every generated token, resulting in low utilization of computational cores and high latency. While various compression methods for KV cache have been proposed to alleviate this issue, they suffer from degradation in generation quality. We introduce TriForce, a hierarchical speculative decoding system that is scalable for long sequence generation. This approach leverages the original model weights and dynamic sparse KV cache via retrieval as a draft model, which serves as an intermediate layer in the hierarchy and is further speculated by a smaller model to reduce its drafting latency. TriForce not only facilitates impressive speedups for Llama2-7B-128K, achieving up to 2.31$\times$ on an A100 GPU but also showcases scalability in handling even longer contexts. For the offloading setting on two RTX 4090 GPUs, TriForce achieves 0.108s/token—only half as slow as the auto-regressive baseline on an A100, which attains 7.78$\times$ on our optimized offloading system. Additionally, TriForce performs 4.86$\times$ than DeepSpeed-Zero-Inference on a single RTX 4090 GPU. TriForce’s robustness is highlighted by its consistently outstanding performance across various temperatures. The code is available at https://github.com/Infini-AI-Lab/TriForce.
@@ 1 Introduction
Large language models (LLMs) with long-context capability, such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and LWM (Liu et al., 2024a) continue to emerge and gain proficient application in scenarios including chatbots, vision generation, and financial analysis (Touvron et al., 2023; Chowdhery et al., 2023; Zhao et al., 2023; Reddy et al., 2024). However, losslessly serving these LLMs efficiently is challenging. Because of the auto-regressive nature of LLMs, the entire key-value (KV) cache, which stores intermediate key-value states from previous contexts to avoid re-computation, together with model parameters will be loaded into GPU SRAM for every token generated, resulting in low utilization of computational cores. In addition to the large volume of model parameters, the memory footprint of KV cache, which grows linearly with sequence length (Pope et al., 2023), is emerging as a new bottleneck for long sequence generation.
@@
Recent methodologies have proposed KV cache eviction strategies (Xiao et al., 2023b; Zhang et al., 2024b; Liu et al., 2024c; Jiang et al., 2023; Ge et al., 2023) to mitigate the substantial memory footprint of KV cache, which selectively discard KV pairs from the cache based on a designed eviction policy, allowing models to generate texts with a limited KV cache budget. However, considering that discarded KV pairs cannot be restored and the difficulty in precisely foreseeing which KV pairs will be crucial for future text generation, they struggle with potential information loss, including hallucination and contextual incoherency (Yang et al., 2024), particularly in long contexts. Such challenges prevent these approaches from boosting speed without sacrificing the performance of models, as illustrated in Figure 1.
@@
Concurrently, speculative decoding, which leverages a lightweight draft model to sequentially predict the next few tokens and let the target model verify the predicted tokens in parallel, is introduced to accelerate LLM inference while provably precisely preserving model output (Leviathan et al., 2023; Chen et al., 2023a; Xia et al., 2024). Nonetheless, deploying it for long sequence generation faces several challenges. First, training draft models to match the context length of target LLMs requires massive computation and it remains questionable whether these small models can achieve the same accuracy with a context length around 1M (Beltagy et al., 2020; Peng et al., 2023; Yan et al., 2024). Second, we found that draft models with existing training-free methods (e.g., KV cache eviction strategies) can result in poor speculating performance. A continuously increasing divergence (Leviathan et al., 2023) is witnessed as the sequence length increases, as shown in Figure 2(a).
@@
Figure 1: Left: TriForce employs retrieval-based drafting and hierarchical speculation to effectively address the dual bottlenecks. It integrates two models and three caches, comprising a draft model, a target model, a StreamingLLM cache for the draft model, alongside a retrieval cache and a full cache for the target model. The process initiates by repeatedly drafting for $\gamma_{1}$ steps, assisting the target model with retrieved partial KV cache in generating over $\gamma_{2}$ tokens, which will be further verified by the target model using full KV cache. Right: Evaluating the Llama2-7B-128K on a needle retrieval task indicates that KV cache eviction-based methods, such as StreamingLLM, require a trade-off between latency and accuracy. In contrast, our TriForce maintains low latency without sacrificing accuracy.
@@
In pursuit of lossless acceleration, we utilize the lossless feature of speculative decoding as the foundation of our system. An ideal speculative decoding algorithm should (i) be training-free, (ii) maintain a high acceptance rate (Leviathan et al., 2023) with long contexts, and (iii) have low-cost drafting. However, two technical challenges need to be addressed to achieve the goal. First, it is not immediately apparent what we can use for low-latency drafting without training a smaller draft model to match the long context length. Second, the key factors for attaining a high acceptance rate with long contexts remain unclear.
@@
Building on these insights, we introduce a hierarchical speculation approach. For a long-context target model (e.g., Llama2-7B-128K (Peng et al., 2023)), we leverage the original model weights but only with a small proportion (e.g., $3\%$) of KV cache as a draft to tackle the bottleneck of KV cache. Hierarchically, the draft model is further speculated by a lightweight model (e.g., Llama-68M) with StreamingLLM cache to address the bottleneck of model weights. We present TriForce, depicted in Figure 1, a scalable and robust speculative decoding system that integrates retrieval-based drafting and hierarchical speculation, optimized for both on-chip and offloading scenarios. Specifically,
  • In Section 4.1, by maintaining the full cache, we gain the flexibility to select KV pairs, allowing us to devise a superior selection method, termed retrieval-based drafting. This strategy retrieves required context information for future needs, which is characterized as lossless, particularly in comparison to eviction-based methods such as StreamingLLM and H2O. We further demonstrated its effectiveness and robustness on different datasets.
@@
Figure 2: (a) A continuously increasing natural divergence (Leviathan et al., 2023) between the draft model with StreamingLLM or Sliding-window with Re-computation and Llama2-7B-128K is witnessed as the sequence length increases, indicating a falling acceptance rate for speculative decoding with longer contexts. Additionally, temperature sensitivity signals a lack of robustness. (b) Compared with model weights, KV cache gradually becomes another bottleneck with long contexts. (c) KV cache occupies most of the memory as the context length increases.
Empirically, in Section 5, we perform extensive experiments and ablation studies to demonstrate the effectiveness of TriForce. We show that TriForce achieves up to 2.31$\times$ speedup for Llama2-7B-128K on a single A100 GPU on top of Hugging Face (Wolf et al., 2019) with CUDA graphs (NVIDIA & Fitzek, 2020). For the offloading setting, TriForce attains an impressive 7.78$\times$ on two RTX 4090 GPUs, reaching 0.108 s/token—only half as slow as the auto-regressive baseline on an A100. TriForce can efficiently serve a Llama2-13B with 128K contexts with 0.226s/token, which is 7.94$\times$ faster on our optimized offloading system. On a single RTX 4090 GPU, TriForce is 4.86$\times$ faster than DeepSpeed-Zero-Inference (Aminabadi et al., 2022)  [^1] The official implementation of DeepSpeed-ZeRO-Inference (Aminabadi et al., 2022) with KV cache offloading currently only supports a single GPU, which computes attention on CPU. Our offloading system transfers KV cache from CPU to GPU, benefiting from Tensor Parallelism.. Further, we show that: (i) TriForce has a theoretical 13.1$\times$ upper bound, demonstrating exceptional scalability when dealing with long contexts; (ii) TriForce is robust across various temperature settings, maintaining an acceptance rate above 0.9 even for a temperature of 1.0; and (iii) TriForce’s ability to efficiently process large batches, achieving a 1.9$\times$ speedup for a batch size of six with 19K contexts per sample.
@@ 2.1 Speculative Decoding
Speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023a; Kim et al., 2024; Zhang et al., 2023; Santilli et al., 2023; Hooper et al., 2023) is featured by accelerating LLM decoding while precisely maintaining the model’s output distribution. As the speed of the auto-regressive decoding process is mainly bound by the time for loading model weights and KV cache to GPU SRAM, speculative decoding leverages the observation that generating one token takes the same time as processing tens of tokens in parallel. Tree-based speculation methods are proposed to fully utilize the speculation budget (Fu et al., 2024; Li et al., 2024). Instead of making one prediction for the next token, tree-based methods leverage multiple candidates to boost the acceptance rate so that more tokens can get accepted (Miao et al., 2023; Sun et al., 2024; Chen et al., 2024). Staged speculation techniques (Spector & Re, 2023; Chen et al., 2023b) have been suggested to further accelerate inference by using a cascade of draft models, and hierarchical speculation shares similarities with these approaches. However, our method focuses on long sequence generation tasks, which presents unique challenges. We use different drafting methods for each speculation phase to address two distinct bottlenecks in long-context scenarios, instead of focusing on model weights at every stage for acceleration. Meanwhile, self-speculation approaches such as Medusa (Cai et al., 2024; Ankner et al., 2024), which are orthogonal to our method, require training efforts and can be integrated into our intermediate draft model.
@@ 2.2 KV Cache Eviction Strategies
StreamingLLM (Xiao et al., 2023b) addresses the limitations of window attention and sliding window with re-computation by presenting a straightforward yet effective method that allows LLMs to handle infinitely long text sequences without fine-tuning. StreamingLLM stabilizes the performance by retaining critical attention sink tokens together with recent KV for attention computation. By prioritizing sink tokens, StreamingLLM ensures the attention score distribution remains stable, promoting consistent language modeling for long texts.
H2O (Zhang et al., 2024b) introduces a greedy but low-cost approach to processing infinite-length input streams, inspired by a simplified version of the heavy-hitters (H2) eviction policy. This method dynamically updates the KV cache based on the cumulative attention scores, systematically removing the least critical KV to maintain a fixed cache size. By leveraging a greedy algorithm based on local statistics, H2O effectively selects which KV pairs to preserve in the cache, ensuring efficient inference without compromising quality.
However, it is important to recognize that these techniques do not increase the context window size (Zhang et al., 2024a; Jin et al., 2024; Ge et al., 2023; Jiang et al., 2023). They focus on retaining only the most recent tokens along with either attention sinks or heavy-hitters, while discarding other tokens. These approaches limit the model to processing based on their designed eviction policies and recent tokens. Consequently, they might not be directly applicable to tasks that demand comprehensive, long-context understanding.
@@ 2.3 KV Cache Quantization
Several approaches to KV cache quantization have been introduced to enhance the efficiency of inference for long sequence generation, aiming to maintain generation quality while reducing the memory consumption (Xiao et al., 2023a; Hooper et al., 2024; Sheng et al., 2023; Liu et al., 2023a; Zirui Liu et al., 2023; Yue et al., 2024). Quantization methods focus on compressing the bit width of KV cache activations, which is orthogonal to our approach.
@@ 3.1 Leveraging Attention Sparsity for Speculative Decoding
The phenomenon of attention sparsity in pre-trained LLMs has been discovered by numerous studies (Zhang et al., 2024b; Xiao et al., 2023b; Liu et al., 2023b; 2024c). In our study, we conduct zero-shot inference on the PG-19 test set (Rae et al., 2019) with Llama2-7B-128K model. By visualizing the sparsity across different attention heads, demonstrated in Figure 3(a), we observe that with a context length of 120K, it is possible to recover over 96% of the attention score with merely 4K tokens across almost all layers.
@@ 4.1 Retrieval-based Drafting
In scenarios requiring long-term contextual dependencies, methods like StreamingLLM and H2O underperform due to their cache updating strategies, which are ineffective at accurately retrieving detailed contextual information because they inevitably and irrecoverably discard KV pairs. In our experiment, we challenge StreamingLLM and H2O with a needle retrieval task (Liu et al., 2024b; Peng et al., 2023; Liu et al., 2024a). As detailed in Table 1, there is a notable drop in their acceptance rates compared to their performance on the PG-19 dataset, highlighting their limitations. Essentially, StreamingLLM and H2O operate on a lossy principle, as evicted tokens are permanently discarded, making them a poor fit for settings requiring the preservation of full KV cache for the target model.
The necessity of keeping the entire KV cache in our settings allows us to select KV cache more freely (Singhal et al., 2001). This insight leads us to develop a more effective selection policy for lossless approximations. In our approach, demonstrated in Figure 4, KV cache is segmented into small chunks. During the retrieval phase, we calculate the attention between a given query and the average key cache within each chunk. This method effectively highlights the most relevant chunks, enabling us to gather KV cache with a fixed budget based on the scores. As illustrated in Table 1, retrieval-based method excels by actively identifying the most crucial information for the task rather than relying on passive and time-based cache management methods. By focusing on relevance over recency, retrieval-based policy demonstrates its potential to handle contextually dense datasets.
Table 1: Acceptance rates are shown across various tasks, utilizing a 120K context and a 4K budget, while bypassing the initial two layers. There is a notable drop in StreamingLLM and H2O’s results on needle retrieval. For reference, Top-K is the theoretical upper bound.
@@ 4.2 Hierarchical Speculation
While addressing the KV cache bottleneck enhances efficiency, the requirement to load whole model weights for drafting reintroduces latency, shifting the bottleneck to model weights again. To tackle this challenge, we implement a hierarchical system, as illustrated in Figure 1. This system employs a secondary, lightweight model with StreamingLLM cache to perform initial speculations for the target model with retrieval-based cache (which serves as a draft model for the target model with full KV cache). By establishing this sequential speculation hierarchy, we effectively reduce drafting latency and accelerate overall inference.
Correctness: The original output distribution is preserved during the final speculation phase, which is identical to the standard speculative decoding algorithm (Leviathan et al., 2023; Chen et al., 2023a), and the proof is trivial.
4.3 Algorithm
TriForce is devised to exploit the bottlenecks associated with both model weights and KV cache to enhance the inference speed of LLMs for long sequence generation. We present the pseudocode for the TriForce in Algorithm 1. It starts by prefilling the target model $M_{p}$ with full cache $C_{p}$ and draft model $M_{q}$ with StreamingLLM cache $C_{q}$ using a given input prefix, and then constructs the retrieval cache $C_{r}$. The initialization and update mechanism for the retrieval cache $C_{r}$ is guided by the insights of contextual locality discussed in Section 3.2. We first construct $C_{r}$ using the last token of the prefix, arranging tokens by the descending order of importance. In subsequent inferences, we overwrite tokens with the least importance, maintaining the relevance and utility of the cache. A reconstruction of $C_{r}$ is triggered either when the rolling average acceptance rate drops below a threshold or at a designed stride.
The inference progresses iteratively until it reaches the target sequence length $T$. After each iteration, cache $C_{r}$ and $C_{q}$ are updated to prepare for the subsequent speculation phase. Each iteration encompasses two speculations: initially, $M_{q}$ utilizes $C_{q}$ to predict $M_{p}$ with $C_{r}$ for $\gamma_{1}$ steps until $n\geq\gamma_{2}$. Subsequently, these $n$ tokens are self-verified (Zhang et al., 2023) by $M_{p}$ with $C_{p}$. This process constructs a hierarchy: the first layer of hierarchy employs a smaller, faster model $M_{q}$ with local context $C_{q}$ to speculate the large model $M_{p}$ with partial but high-quality global context $C_{r}$, addressing the model weights bottleneck. The second layer utilizes model $M_{p}$ with retrieval cache for self-speculation, overcoming the bottleneck caused by KV cache. This hierarchical speculation algorithm boosts efficiency by effectively addressing both bottlenecks. System implementation is detailed in Appendix A.